# Vulnerability of LLMs to Vertically Aligned Text Manipulations

### v1

@@ Abstract
Text classification involves categorizing a given text, such as determining its sentiment or identifying harmful content. With the advancement of large language models (LLMs), these models have become highly effective at performing text classification tasks. However, they still show vulnerabilities to variations in text formatting. Recent research demonstrates that modifying input formats, such as vertically aligning words for encoder-based models, can substantially lower accuracy in text classification tasks. While easily understood by humans, these inputs can significantly mislead models, posing a potential risk of bypassing detection in real-world scenarios involving harmful or sensitive information. With the expanding application of LLMs, a crucial question arises: Do decoder-based LLMs exhibit similar vulnerabilities to vertically formatted text input? In this paper, we investigate the impact of vertical text input on the performance of various LLMs across multiple text classification datasets and analyze the underlying causes. Our findings are as follows: (i) Vertical text input significantly degrades the accuracy of LLMs in text classification tasks. (ii) Chain of Thought (CoT) reasoning does not help LLMs recognize vertical input or mitigate its vulnerability, but few-shot learning with careful analysis does. (iii) We explore the underlying cause of the vulnerability by analyzing the inherent issues in tokenization and attention matrices.
@@ 3.2 Word Transformation
Model SST-2 CoLA QNLI Rotten Tomatoes Jigsaw Toxicity Original Vertical Original Vertical Original Vertical Original Vertical Original Vertical Closed-Source Models GPT-3.5 93.00 65.00 $(\downarrow\textbf{28.00})$ 80.00 47.00 $(\downarrow\textbf{33.00})$ 85.00 69.00 $(\downarrow\textbf{16.00})$ 92.00 57.00 $(\downarrow\textbf{35.00})$ 85.00 62.00 $(\downarrow\textbf{23.00})$ GPT-4 96.00 67.00 $(\downarrow\textbf{29.00})$ 90.00 49.00 $(\downarrow\textbf{41.00})$ 89.00 71.00 $(\downarrow\textbf{18.00})$ 93.00 64.00 $(\downarrow\textbf{29.00})$ 89.00 58.00 $(\downarrow\textbf{31.00})$ GPT-4o-mini 95.00 66.00 $(\downarrow\textbf{29.00})$ 89.00 50.00 $(\downarrow\textbf{39.00})$ 90.00 71.00 $(\downarrow\textbf{19.00})$ 91.00 61.00 $(\downarrow\textbf{30.00})$ 85.00 57.00 $(\downarrow\textbf{28.00})$ GPT-4o 95.00 68.00 $(\downarrow\textbf{27.00})$ 87.00 47.00 $(\downarrow\textbf{40.00})$ 90.00 70.00 $(\downarrow\textbf{20.00})$ 90.00 65.00 $(\downarrow\textbf{25.00})$ 91.00 60.00 $(\downarrow\textbf{31.00})$ Open-Source Models Llama3-8B 89.00 61.00 $(\downarrow\textbf{28.00})$ 75.00 50.00 $(\downarrow\textbf{25.00})$ 83.00 62.00 $(\downarrow\textbf{21.00})$ 86.00 42.00 $(\downarrow\textbf{44.00})$ 88.00 58.00 $(\downarrow\textbf{30.00})$ Llama3-70B 96.00 67.00 $(\downarrow\textbf{29.00})$ 85.00 48.00 $(\downarrow\textbf{37.00})$ 86.00 63.00 $(\downarrow\textbf{23.00})$ 91.00 46.00 $(\downarrow\textbf{45.00})$ 88.00 58.00 $(\downarrow\textbf{30.00})$ Llama3.1-8B 93.00 51.00 $(\downarrow\textbf{42.00})$ 80.00 49.00 $(\downarrow\textbf{31.00})$ 83.00 59.00 $(\downarrow\textbf{24.00})$ 89.00 54.00 $(\downarrow\textbf{45.00})$ 80.00 64.00 $(\downarrow\textbf{16.00})$ Llama3.1-70B 96.00 66.00 $(\downarrow\textbf{30.00})$ 84.00 50.00 $(\downarrow\textbf{34.00})$ 84.00 66.00 $(\downarrow\textbf{18.00})$ 92.00 63.00 $(\downarrow\textbf{29.00})$ 87.00 62.00 $(\downarrow\textbf{25.00})$ Gemma2-9B 88.00 60.00 $(\downarrow\textbf{28.00})$ 83.00 51.00 $(\downarrow\textbf{32.00})$ 78.00 60.00 $(\downarrow\textbf{18.00})$ 86.00 58.00 $(\downarrow\textbf{28.00})$ 85.00 53.00 $(\downarrow\textbf{32.00})$ Gemma2-27B 94.00 58.00 $(\downarrow\textbf{36.00})$ 87.00 50.00 $(\downarrow\textbf{37.00})$ 83.00 64.00 $(\downarrow\textbf{19.00})$ 89.00 54.00 $(\downarrow\textbf{35.00})$ 88.00 55.00 $(\downarrow\textbf{33.00})$ Qwen1.5-72B 95.00 63.00 $(\downarrow\textbf{32.00})$ 81.00 52.00 $(\downarrow\textbf{29.00})$ 85.00 66.00 $(\downarrow\textbf{19.00})$ 93.00 57.00 $(\downarrow\textbf{36.00})$ 88.00 63.00 $(\downarrow\textbf{25.00})$ Qwen2-72B 96.00 60.00 $(\downarrow\textbf{36.00})$ 84.00 50.00 $(\downarrow\textbf{34.00})$ 88.00 62.00 $(\downarrow\textbf{26.00})$ 93.00 59.00 $(\downarrow\textbf{34.00})$ 91.00 59.00 $(\downarrow\textbf{32.00})$
@@ 4.2 Large Language Models
Closed-Source Models. We utilize four different models developed by OpenAI, including the latest GPT-4o and GPT-4o-mini.
@@ 4.3 Metric
4.4 Implementation Details
For the other open-source LLMs, we use llamaapil [^3] https://www.llama-api.com/ and groq [^4] https://groq.com/ platforms for model inference, applying the same parameter settings as those used for the GPT series models.
@@ 4.5 Experiment Results
We utilize four GPT series LLMs along with eight widely recognized open-source LLMs on five text classification datasets. For the SST-2, QNLI, Rotten Tomatoes, and Jigsaw Toxicity datasets, the GPT-4o-mini model would select four words for vertical input. In contrast, for the CoLA dataset, it selects two words. In the label prediction phase, we utilize straightforward and precise prompts for zero-shot prediction, enabling the model to generate predicted labels. We compare the classification accuracy before and after modifying the input format, with the results presented in Table 1.
@@
Specifically, for the CoLA dataset, which determines whether a sentence is grammatically correct, the model’s classification accuracy drops by nearly 40 percentage points when two key words are input vertically. Given that the random prediction accuracy for a binary classification problem is 50%, this indicates that the model essentially loses its classification ability.

### v3

@@ Abstract
Vertical text input is commonly encountered in various real-world applications, such as mathematical computations and word-based Sudoku puzzles. While current large language models (LLMs) have excelled in natural language tasks, they remain vulnerable to variations in text formatting. Recent research demonstrates that modifying input formats, such as vertically aligning words for encoder-based models, can substantially lower accuracy in text classification tasks. While easily understood by humans, these inputs can significantly mislead models, posing a potential risk of bypassing detection in real-world scenarios involving harmful or sensitive information. With the expanding application of LLMs, a crucial question arises: Do decoder-based LLMs exhibit similar vulnerabilities to vertically formatted text input? In this paper, we investigate the impact of vertical text input on the performance of various LLMs across multiple text classification datasets and analyze the underlying causes. Our findings are as follows: (i) Vertical text input significantly degrades the accuracy of LLMs in text classification tasks. (ii) Chain of Thought (CoT) reasoning does not help LLMs recognize vertical input or mitigate its vulnerability, but few-shot learning with careful analysis does. (iii) We explore the underlying cause of the vulnerability by analyzing the inherent issues in tokenization and attention matrices.
@@ 3.2 Word Transformation
Model SST-2 CoLA QNLI Rotten Tomatoes Jigsaw Toxicity Original Vertical Original Vertical Original Vertical Original Vertical Original Vertical Closed-Source Models GPT-3.5 93.00 65.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{28.00}})$ 80.00 47.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{33.00}})$ 85.00 69.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{16.00}})$ 92.00 57.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{35.00}})$ 85.00 62.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{23.00}})$ GPT-4 96.00 67.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{29.00}})$ 90.00 49.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{41.00}})$ 89.00 71.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{18.00}})$ 93.00 64.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{29.00}})$ 89.00 58.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{31.00}})$ GPT-4o-mini 95.00 66.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{29.00}})$ 89.00 50.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{39.00}})$ 90.00 71.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{19.00}})$ 91.00 61.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{30.00}})$ 85.00 57.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{28.00}})$ GPT-4o 95.00 68.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{27.00}})$ 87.00 47.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{40.00}})$ 90.00 70.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{20.00}})$ 90.00 65.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{25.00}})$ 91.00 60.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{31.00}})$ Open-Source Models Llama3-8B 89.00 61.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{28.00}})$ 75.00 50.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{25.00}})$ 83.00 62.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{21.00}})$ 86.00 42.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{44.00}})$ 88.00 58.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{30.00}})$ Llama3-70B 96.00 67.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{29.00}})$ 85.00 48.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{37.00}})$ 86.00 63.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{23.00}})$ 91.00 46.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{45.00}})$ 88.00 58.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{30.00}})$ Llama3.1-8B 93.00 51.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{42.00}})$ 80.00 49.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{31.00}})$ 83.00 59.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{24.00}})$ 89.00 54.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{45.00}})$ 80.00 64.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{16.00}})$ Llama3.1-70B 96.00 66.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{30.00}})$ 84.00 50.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{34.00}})$ 84.00 66.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{18.00}})$ 92.00 63.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{29.00}})$ 87.00 62.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{25.00}})$ Gemma2-9B 88.00 60.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{28.00}})$ 83.00 51.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{32.00}})$ 78.00 60.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{18.00}})$ 86.00 58.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{28.00}})$ 85.00 53.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{32.00}})$ Gemma2-27B 94.00 58.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{36.00}})$ 87.00 50.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{37.00}})$ 83.00 64.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{19.00}})$ 89.00 54.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{35.00}})$ 88.00 55.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{33.00}})$ Qwen1.5-72B 95.00 63.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{32.00}})$ 81.00 52.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{29.00}})$ 85.00 66.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{19.00}})$ 93.00 57.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{36.00}})$ 88.00 63.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{25.00}})$ Qwen2-72B 96.00 60.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{36.00}})$ 84.00 50.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{34.00}})$ 88.00 62.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{26.00}})$ 93.00 59.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{34.00}})$ 91.00 59.00 $(\downarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{32.00}})$
@@ 4.2 Large Language Models
Closed-Source Models. We utilize four different models developed by OpenAI, including the latest powerful models GPT-4o and GPT-4o-mini.
@@ 4.4 Experiment Results
We utilize four GPT series LLMs along with eight widely recognized open-source LLMs on five text classification datasets. For the SST-2, QNLI, Rotten Tomatoes, and Jigsaw Toxicity datasets, the GPT-4o-mini model would select four words for vertical input. In contrast, for the CoLA dataset, it selects two words.
In the label prediction phase, we utilize straightforward and precise prompts for zero-shot prediction, enabling the model to generate predicted labels. We compare the classification accuracy before and after modifying the input format, with the results presented in Table 1. The detailed implementation of different LLMs is shown in Appendix B.
@@
Specifically, for the CoLA dataset, the model’s classification accuracy drops by nearly 40 percentage points when two key words are input vertically. Given that the random prediction accuracy for a binary classification problem is 50%, this indicates that the model essentially loses its classification ability.
@@
For the other three datasets, the decrease in accuracy after implementing vertical input for different models mostly ranges from 25 to 40 percentage points, which is a significant drop and almost indicates that current LLMs nearly lose their natural language understanding ability when encountering vertical input, even when the manipulation involves only a few key words.